OpenStudy (anonymous):

Felix purchased a boat for $9,755. He made a down payment of $875. He applied for a six-year installment loan with an interest rate of 10.4% in the amount of $8,880. What is the total cost of the boat after six years?

OpenStudy (anonymous):

\[[1-(1.104^{\neg6})]\div0.104 \approx 4.3047\] \[8880\div4.3047 \approx 2063\] \[2063\times6 = 12378\] \[12378+875=13253\] Total cost = $13,253 Note: Interest cost = 12378 - 8880 = $3,498